- Current status as of 1/22: IR and Green reached X end PD.
1/22: Surface inspection for PR mirrors. Cleaning will be done after 1/31.
1/23: Surface inspection for SR mirrors.
1/24-25: Cleaning HR of SR3, SRM.
1/23: BS is aligned to Y-end PD by IR.
1/27: ITMY is aligned to REFL by IR.
1/27: ITMX is aligned to REFL by IR.
1/27: Beam position of IR at SR3 is confirmed how far from the center.
1/28-29: SR3 is aligned to SR2 center by IR.
1/28: Alignment of SR2 -> SRM by IR.
1/29: Green beam is injected from SR2 AR surface to Y end PD.
- 1/31: Removing bellows at PR and SR area.
- 1/31-2/1: Cleaning HR of PR3 and HR, AR of SR2
2/4-8: SRM -> OMMT by IR -> OMMT -> OSTM by IR, OSTM -> OMC by IR
2/4-8: Alignment of ETMY -> POS by Green
- 2/12-15: Closing door for central area, pumping down
- needs drift monitor for VIS
Report of work (*:responsible person)
- Type B: Height issues are being fixed now. RT model for SRs should be separated to upper slow one and lower fast one like BS.
- PRM: Upgrading oplev.
- IOO: High power beam shutter mirror was dropped. It will be fixed today.
- ETMY: Stepping motor was not working.
- AOS: heatsink for ETMX.
Work plan (*:responsible person)
- VAC: removing bellows at PR3 and SR2.
- MIR: inspection and cleaning PR3 and SR2 by tomorrow.